Please help me solve this problem:
The length of a rectangle is 20ft. and the diagonal is 35ft. What is the perimeter of the rectangle

Answer by Euler(11)

Okay, so here, you'll apply Pythagoras Theorem:
20^2+x^2=35^2
35^2-20^2=x^2
825=x^2
Taking square root on both sides, we get:
28.723=x
My answer's rounded off to 3 dp. :)